OpenStudy (anonymous):

an arrow is fired directly horizontal off a cliff that is 10 meters tall with a velocity of 65.5m/s. A. How long is the arrow in the air? B. What is the range of the arrow?

OpenStudy (anonymous):

\[\Delta y= -10 = -\frac{ g t^2 }{ 2 }\]

OpenStudy (anonymous):

find t and use in \[\Delta x = 65.5t\]

OpenStudy (anonymous):

\[\Delta y = \frac{ -1 }{ 2 } (g) (t^2)\]

OpenStudy (anonymous):

\[\Delta x = Vx(t)\]

OpenStudy (anonymous):

\[g=9.8\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

which one is x and which one is y?

OpenStudy (shamim):

i can solve it

OpenStudy (shamim):

r u rhere

OpenStudy (shamim):

h=0.5gt*t or, 10=0.5*9.8*t*t or, t=1.4 sec

OpenStudy (shamim):

range R=65.5*1.4=91.7m

OpenStudy (shamim):

if u need explanation then ask me when u r online